OpenStudy (anonymous):

Solve:(-2x-5)+7x>-15

OpenStudy (anonymous):

Someone plz help me

OpenStudy (jack1):

\[(-2x-5)+7x>-15\] \[-2x-5+7x>-15\] \[7x-2x-5>-15\] \[5x-5>-15\] \[5x>-15+5\] \[5x>-10\]... can you solve from here @Punkchick ?

OpenStudy (anonymous):

no

OpenStudy (jack1):

sók, next step is to divide both sides by 5 so 5x/5 = ? and -10/5 = ? once you've done that, you'll have your final answer

OpenStudy (anonymous):

would it be x<-2??

OpenStudy (jack1):

almost perfect, the symbol doesnt change direction though so x>-2

OpenStudy (anonymous):

Can u help me on a couple more jack 1???

OpenStudy (jack1):

sweet, shoot

OpenStudy (anonymous):

Solve: 2 (b-8)>12

OpenStudy (judygreeneyes):

you got this one already ;) in the other post

OpenStudy (anonymous):

No I don't think so.

OpenStudy (jack1):

\[2 (b-8)>12\] expand out the brackets \[2b-16>12\] now add 16 to both sides \[2b>12+16\] now isolate the b by dividing both sides by 2 \[2b>28\] =...?

OpenStudy (anonymous):

14

OpenStudy (jack1):

spot on, so b>14

OpenStudy (anonymous):

Solve: 5x-6<9x+30

OpenStudy (jack1):

\[5x-6<9x+30\] ok i'll tell u the steps, and you do this one, k? \[5x-6<9x+30\] first, add 6 to both sides next, minus (9x) from both sides now you should have all the x numbers on one side, and all the plain numbers on the other side so finally, divide both sides by whatever number is left in front of the x ... = final answer now show me ur steps

OpenStudy (anonymous):

???\

OpenStudy (anonymous):

13,5

OpenStudy (anonymous):

Am I I right???

OpenStudy (jack1):

mmm... could u show me the steps u went through, line by line, coz i got a diffrnt answer

OpenStudy (anonymous):

I really need to know the answer

OpenStudy (anonymous):

Could u plz tell me it

OpenStudy (anonymous):

Am I right???

OpenStudy (jack1):

\[5x−6<9x+30\] first, add 6 to both sides \[5x−6+6<9x+30+6\] next, minus (9x) from both sides \[5x−6+6-9x<9x+30+6-9x\] now the -6+6 = 0 on the left hand side and the 9x - 9x= 0 on the right hand side so... \[5x-9x<+30+6\] now you should have all the x numbers on one side, and all the plain numbers on the other side 5x - 9x = ? and 30 + 6 = ? so finally, divide both sides by whatever number is left in front of the x ... = final answer

OpenStudy (anonymous):

32

OpenStudy (jack1):

5x - 9x = -4x amd 30 + 6 = 36 so -4x<36 so times both sides by -1 means you flip the direction of the sign 4x>-36 x > -9